OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle.

OpenStudy (anonymous):

b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

what is c.)? Thank you for the above help

OpenStudy (nurali):

sin(30)=0.5000 cos(30)=0.8660 tan(30)=0.5773

OpenStudy (dumbcow):

sin = opposite/hypotenuse = 41/82 cos = adjacent/hypotenuse = 41sqrt3 / 82 tan = opposite/adjacent = 41/41sqrt3

OpenStudy (anonymous):

thank you!!!

OpenStudy (nurali):

My Pleasure